OpenStudy (anonymous):

Use the rules of exponents to evaluate or simplify. Write without negative exponents. (bm) -3

OpenStudy (debbieg):

\(\large (bm)^{-3}\) Use the following rules: \(\large (ab)^{n}=a^n\cdot b^n\) \(\large a^{-1}=\dfrac{1}{a}\) \(\large a^{-n}=\dfrac{1}{a^n}\)